OpenStudy (anonymous):

5(1.15)^t = 8(1.07)^t solve for t exactly if possible

OpenStudy (anonymous):

If you could list your process as well that would be greatly appreciated

OpenStudy (anonymous):

First i would get like terms on one side 5/8=(1.07/1.15)^t Now to get rid of the exponent i would use the either log function (ln or log) ln(5/8)=tln(1.07/1.15) Then (ln(5/8))/(ln(1.07/1.15))=t

OpenStudy (anonymous):

Thanks!

OpenStudy (campbell_st):

you need to take the log of both sides and use some log laws use base e logs \[\ln(5\times(1.15)^t) = \ln(5 \times (1.07)^t)\] this can be simplified to \[\ln(5) + t \times \ln(1.15) = \ln(8) + t \times \ln(1.07)\] or \[t ( \ln(1.15) - \ln(1.07)) = \ln(8) - \ln(5)\] that should help you solve for t